# Approximately Dominating Sets in Elections

### v1

@@ 1.3 Concurrent Work
Independent to our work, a new preprint by Bourneuf, Charbit, and Thomassé [BCT25] also shows that elections have approximately dominating sets of constant size, specifically $(\frac{1}{2}-\varepsilon)$-dominating sets of size $O(\frac{1}{\varepsilon^{2}}\log\frac{1}{\varepsilon})$ — the same as the first part of Theorem 1, but with an additional $O(\log\frac{1}{\varepsilon})$ factor. As a point of comparison, [BCT25] guarantee $0.49$-dominating sets of size 5,227,032, a bound that they highlight as far from reality. On the other hand, Theorem 1 would give roughly $\frac{\pi}{8\times 0.01^{2}}\approx 3,927$, which is still quite large but at least plausible for applications like NSF grants or certain academic conferences where the number of acceptances is in the thousands. For these practical applications, the more general “average-case” guarantee of Theorem 1 may also be appealing — one can ensure that any rejected grant or paper is not much better than the (weighted) average of those that are accepted.
Another intriguing difference is in the techniques that are used to obtain these results. [BCT25]’s approach can be thought of as extending [ABK+06], which argues that majority graphs derived from elections of quality $\varepsilon$ must have bounded VC-dimension. (Note that [ABK+06] and [BCT25] get the same bound.) [BCT25] adapt majority graphs to allow two kinds of edges (differentiating large and small margins between candidates) and show that an appropriate generalization of the VC-dimension is still bounded using the dense neighborhood lemma, a tool whose versatility they showcase through a wide range of applications.
The differing approaches between [BCT25] and our work offer distinct advantages. Our approach is somewhat more direct, and gives a sharper quantitative guarantee that also extends to distributions over candidates with small support. On the other hand, [BCT25] show that their techniques can be applied to a broader suite of combinatorial structures beyond elections. It would be exciting to understand whether the best of both worlds is possible, and more generally to explore deeper connections between the two approaches.

### v2

@@ 1.3 Concurrent Work
Independent to our work, a new preprint by Bourneuf, Charbit, and Thomassé [BCT25] also shows that elections have approximately dominating sets of constant size, specifically $(\frac{1}{2}-\varepsilon)$-dominating sets of size $O(\frac{1}{\varepsilon^{2}})$ — the same as the first part of Theorem 1. Intriguingly, our results are proved via very different techniques. [BCT25]’s approach can be thought of as extending [ABK+06], which argues that majority graphs derived from elections of quality $\varepsilon$ must have bounded VC-dimension. [BCT25] adapt majority graphs to allow two kinds of edges (differentiating large and small margins between candidates) and show that an appropriate generalization of the VC-dimension is still bounded using the dense neighborhood lemma, a tool whose versatility they showcase through a wide range of applications.
These differing approaches offer distinct advantages. Our approach is somewhat more direct, and gives a sharper quantitative guarantee that also extends to distributions over candidates with small support. On the other hand, [BCT25] show that their techniques can be applied to a broader suite of combinatorial structures beyond elections. It would be exciting to understand whether the best of both worlds is possible, and more generally to explore deeper connections between the two approaches.